OpenStudy (anonymous):

Over a span of 6.0 seconds, a car changes its speed from 89 km/h to 37 km/h. What is its average acceleration in meters per second squared (m/s2)?

OpenStudy (anonymous):

1km /h= 1000/3600=5/18 m/s i think you can do this now